OpenStudy (anonymous):

Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. The equation is 2x + 3y = 1,200. Help Please!

OpenStudy (anonymous):

Alright. Hi there naturegirl. Would you agree with me that Slope-Intercept forms looks like y = mx + b

OpenStudy (anonymous):

Yes, they do

OpenStudy (anonymous):

Now, give the equation 2x + 3y = 1200 what do we have to do in order to make it look like slope intercept form?

OpenStudy (anonymous):

Do we have to move the 3y to the other side of the equation?

OpenStudy (anonymous):

Well. y = mx + b We need the y on one side and the slope and y-intercept (b) on another. It would be easier to keep the y and move the 2x to the other side. How would we do that?

OpenStudy (anonymous):

I don't really know. I'm sorry

OpenStudy (anonymous):

That okay. So what we want to do is subtract 2x from both sides it cancels on the left and goes to the right.